# Extending Audio Context for Long-Form Understanding in Large Audio-Language Models

### v1

@@ Abstract
Large Audio-Language Models (LALMs) are often constrained by short audio context windows, even when their text backbones support long contexts, limiting long-form audio understanding. Prior work has introduced context-extension methods (e.g. YaRN) on unimodal LLMs, yet their application to LALMs remains unexplored. First, building on RoPE-based context extension, we introduce Partial YaRN, a training-free, audio-only extension method that modifies only audio token positions, leaving text positions intact to preserve the base LLM’s text capabilities. Second, we propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a training-time positional augmentation. VLAT simulates diverse audio lengths during training, enabling generalization to inputs far longer than those seen in training and improving robustness for long-context audio understanding. Our experiments on SALMONN and Qwen2-Audio show that Partial YaRN outperforms the original models across wide range of settings, and VLAT training strategy provides substantial improvement, achieving strong performance on long audio of unseen lengths. [^1] Code is available at: https://github.com/scb-10x/partial-yarn.
@@
Yuatyong Chaichana1,2thanks: Work done while interning at SCB 10X., Pittawat Taveekitworachai122footnotemark: 2, Warit Sirichotedumrong122footnotemark: 2, Potsawee Manakul122footnotemark: 2, Kunat Pipatanakul1thanks: Equal advising. 1SCB 10X, SCBX Group, 2Chulalongkorn University Correspondence: potsawee@scb10x.com
Figure 1: An illustration of Partial YaRN. The top plot displays low (blue) and high (red) frequency dimensions of RoPE. The bottom plot displays RoPE with Partial YaRN applied. In this example, the audio window of the low frequency dimensions are interpolated (stretched) by a factor of 1.75x to accommodate longer audio inputs, while the high frequency dimensions remain unaltered to preserve local distances and high frequency information (Peng et al., 2024). Within the Extended Audio Region, the magnitudes of every dimension is scaled to serve as a proxy for attention temperature scaling. Despite that the scaling can create a minor discontinuity in the signal, information from other dimensions can help the model disambiguate this.
@@ 1 Introduction
The success of Large Language Models (LLMs) has spurred multimodal extensions, notably large audio-language models (LALMs) [^2] Also known by other names such as Speech-aware LLMs. that pair an audio encoder with a text backbone. By aligning audio and text in a shared representation, LALMs can leverage the base LLM’s knowledge for complex audio understanding. However, practical use is constrained by short audio context windows: models such as SALMONN (Tang et al., 2024) and Qwen2-Audio (Chu et al., 2024) are typically trained on audio segments of 30s or less, and thus generalize poorly to longer inputs (Guo et al., 2025).
@@
This potential drawback motivates a more targeted strategy. We introduce “Partial YaRN”, an audio-only context extension method that modifies only the audio tokens’ positional encodings. This design enables a direct comparative study against the whole-context approaches, allowing us to investigate the tradeoff between preserving position encodings of the text modality and maintaining a globally uniform positional space.
Our experiments on both training-free and fine-tuned audio context extensions reveal that extension-based methods outperforms original models. However, the best choice between whole-context and audio-only extensions is often model-dependent without a universally superior solution. This suggests that to build truly robust long-audio-context models, we must address the core problem of generalization during the training process. Therefore, we extend Partial YaRN into a novel fine-tuning technique called “Virtual Longform Audio Training” (VLAT). Acting as a positional augmentation technique, it simulates a wide range of audio lengths during fine-tuning, teaching the model to generalize beyond the lengths present in the training dataset. Through VLAT, we obtain excellent results on longform audio of unseen lengths. Our main contributions are as follows:
@@
  • We propose Virtual Longform Audio Training (VLAT), a training strategy that applies Partial YaRN as positional augmentation to simulate diverse audio lengths during training, improving long-context generalization.
  • We present a comparative study between existing context extension methods designed for unimodal LLMs and our proposed methods for LALMs, analyzing their performance trade-off in both training-free and fine-tuning settings.
@@ 2.1 Rotary Positional Encoding
As Transformer (Vaswani et al., 2017) utilizes attention mechanism (Bahdanau et al., 2014) which is position-invariant, explicit positional information has to be provided for the models to differentiate tokens from different positions. Our study focuses on LALMs that use Rotary Positional Encoding (RoPE) (Su et al., 2024), a widely adopted positional encoding.
@@
$\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{k}_{n}$ $=(\boldsymbol{R}_{m,\boldsymbol{\theta}}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n,\boldsymbol{\theta}}\boldsymbol{k}_{n})$ (1)
$=\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{R}_{n-m,\boldsymbol{\theta}}\boldsymbol{k}_{n}\,$, (2)
@@ 2.2 Context Length Extension
Chen et al. (2023) proposes that instead of extrapolating RoPE further to the region unfamiliar for the trained model, interpolation is done inside the pretrained context window. Specifically, let $L$ be the original context length of the model, and $L^{\prime}$ be the target length to be extended to. Define $s=\frac{L^{\prime}}{L}$ as the extension factor. PI simply downscales the base rotational frequency by the extension factor: $\frac{\theta_{i}}{s}$. For example, if originally each positional step is rotated by $\theta_{i}=10^{\circ}$; to double the context length, the frequency (step size) can simply be halved to $5^{\circ}$. This way, the model will never have to attend to positions outside of its trained window; resulting in a more stable approach.
@@
$\operatorname{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}\left(\frac{\boldsymbol{Q}^{T}\boldsymbol{K}}{t\sqrt{d_{k}}}\right)\boldsymbol{V},$
@@
Ding et al. (2024) propose LongRoPE to account for non-uniformity across RoPE dimensions through a novel evolutionary search algorithm to determine dimension-wise interpolation factors. This method is further improved in LongRoPE2 (Shang et al., 2025).
@@ 2.3 Large Audio-language Models and Their Audio Contexts
One broadly adopted paradigm of large audio-language models (LALMs) is the unified input space (Tang et al., 2024; Chu et al., 2024; Ding et al., 2025; Goel et al., 2025) which involves transforming audio inputs into the textual input space of a base LLM. By sharing the embedding space, the strong text capability and learned knowledge of the base LLM can be leveraged while simultaneously being augmented with audio understanding ability. There also exists other LALM paradigms such as cross attention (Kong et al., 2024; Ghosh et al., 2025) where the audio and text modalities are fused through cross-attention modules. This work focuses on models under the unified-input-space.
As with text context window in LLMs, LALMs also do not generalize well to audio with lengths much longer than the window seen during training (e.g. 7s for Pengi (Deshmukh et al., 2023); 30s for Qwen2-Audio (Chu et al., 2024)). Concurrent to us, Guo et al. (2025) introduces FastLongSpeech, an efficient longform LALM that uses Iterative Fusion to compress long audio representation into compact forms. Our work, in contrast, investigates an orthogonal direction of adapting the model’s positional encodings rather than the audio embeddings, and therefore is generally applicable to the existing models without requiring costly retraining.
@@ 3 Partial YaRN
To extend existing unified input LALMs to longer audio, we hypothesize that LALMs already possess a sufficient general understanding of audio and text, but the bottleneck is their unfamiliarity to audio positions beyond the range seen during audio-text training. [^3] This differs from whole-context extension in unimodal LLMs, where the total sequence length exceeds the pre-trained context window, causing an out-of-distribution problem. Here, the total sequence (audio + text) usually still fits within the original window, making the core challenge one of adapting to an unfamiliar audio length and positions, rather than extrapolating to completely unseen positions. This hypothesis suggests a training-free solution: manipulating the backbone LLM’s positional encoding. The goal is to remap positions of a long audio input into the model’s familiar audio range.
First, we study the application of YaRN (Peng et al., 2024) to LALMs, as an example of whole-context extension methods. Since the audio window is a part of the total context window, performing whole-context extension extends the audio window as a byproduct. However, this alters positional information for all tokens, including text, risking the degradation of the base LLM, which was pretrained solely in text. Motivated by the aforementioned hypothesis, we propose Partial YaRN, an audio-only extension method designed to exclusively stretch the audio window. By leaving the text’s positional encodings unaltered, this approach aims to extend [^4] The terms extend and stretch are used interchangeably. audio context while preserving the base model’s text capability.
@@ 3.1 Methodology
Inspired by PI (Chen et al., 2023) and YaRN (Peng et al., 2024) for whole-context extension in unimodal LLMs, we adapt the interpolation technique to LALMs by applying it exclusively to the audio region of the base language models.
@@
where $\oplus$ denotes concatenation in the temporal axis. Note that while the final position index is denoted by $L$, the total token capacity now increases to $L+L^{\prime}_{\text{audio}}-L_{\text{audio}}$ to account for the stretched audio window of a higher token frequency.
@@
We provide empirical validation and a more in-depth discussion of the two-group partitioning in Section 7.2 and Appendix B respectively. See Figure 1 for an example depiction of Partial YaRN.
@@ 4.1 Data
We source audio from the English subset of YODAS2 (Li et al., 2023b). We segment audio samples into non-overlapping segments of 1, 2, 5, and 10 minutes. For each segment, we generate five multiple-choice question-answering (MCQA) pairs using Gemini 2.0 Flash (Team et al., 2023), with four choices each. We design the generation prompt to ensure that each question focuses on different portions of the audio, and that they collectively cover the entire audio segment. The test set of each audio duration (1, 2, 5, and 10 minutes) has 750 QA pairs.
@@ 4.3 Baseline Methods
These three baselines require no tuning as Vanilla and Whole PI don’t utilize any hyperparameter, and YaRN has predefined cutoffs and a closed-form formula for obtaining temperature.
@@
We also report performance of two proprietary models capable of long audio context: GPT-4o (OpenAI, 2024) and Gemini 2.0 Flash (Team et al., 2023), in order to provide a broader perspective of LALMs and to validate the quality of our dataset [^5] The model versions are gpt-4o-mini-2024-07-18 and gemini-2.0-flash-001 respectively. Note that Gemini was also used for generating our YODAS2-MCQA dataset..
@@ 5 Main Results
In this section, we comparatively study the effect of whole-context and audio-only context extensions on both the training-free and fine-tuning settings for audio segments up to 10mins.
@@ 5.1 Training-free Audio Context Extension
First, we extend the audio context length without training. We report the results in Table 4.1, showing that: no single method—whole-context or audio-only—is universally superior. Instead, the performance varies between the different models and the degree of extension. More importantly, we observe that both models generally retain consistent performance up to 2mins, suggesting that their innate audio context lengths are longer than just 30s, possibly a result from multi-audio training.
@@
As previously mentioned in the beginning of this section that we can observe both model’s innate audio contexts to be around 2mins instead of just 30s, here we experiment with extending from this observed audio context of 2mins to the target lengths. For both SALMONN and Qwen2-Audio, performance on the long audio improves substantially when the interpolation is anchored from a 2mins context instead of the original 30s. Notably on Qwen2-Audio, this strategy boosts Partial YaRN’s performance at 10mins from 28.53% to a 48.00%, an absolute improvement of almost 20%, yielding a final performance 26% higher than the Vanilla baseline. This finding suggests that determining the extension by observing the vanilla performance may be a more critical factor than the choice between whole-context or audio-only methods, especially on large extension ratios.
@@ 5.2 Fine-tuning for Audio Context Extension
The results are presented in Table 3, showing that fine-tuning with a context extension method dramatically outperforms the vanilla fine-tuning baseline, especially at longer audio lengths. In the qkvo setting, both Whole YaRN (83.47%) and Partial PI (83.07%) achieve an absolute performance gain of around 19% over the Vanilla baseline (64.93%) at 10mins. This highlights the benefit of integrating a context extension method into the fine-tuning process for more better context extension.
Comparing the whole-context and audio-only approaches, we find that their performance is generally competitive. While Whole YaRN triumps when only adapting the query projection weights, suggesting it’s ease of adaptation. Their performance converges under standard LoRA practice which adapts qkvo. This result confirms that both whole-context and audio-only methods are highly effective strategies for fine-tuning.
@@
Table 3: Finetuned audio context extension performance (Accuracy). We finetune Qwen2-Audio on YODAS2-MCQA, under two LoRA settings: adapting only q, and adapting qkvo. Each model is finetuned and evaluated using the same extension method.
@@ 6 Virtual Longform Audio Training: Generalizing Beyond the Training Data
Our previous results show that while the training-free methods are beneficial, their effectiveness is often model and audio-length dependent. Based on this, we propose to fine-tune LALMs with Partial YaRN repurposed as a positional augmentation technique we call "Virtual Longform Audio Training" (VLAT).

### v2

@@ Abstract
Large Audio-Language Models (LALMs) are often constrained by short audio context windows, even when their text backbones support long contexts, limiting long-form audio understanding. Prior work has introduced context-extension methods (e.g. YaRN) on unimodal LLMs, yet their application to LALMs remains unexplored. First, building on RoPE-based context extension, we introduce Partial YaRN, a training-free, modality-decoupled extension method that modifies only audio token positions, leaving text positions intact to preserve the base LLM’s text capabilities. Second, we propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a training-time positional augmentation. VLAT simulates diverse audio lengths during training, enabling generalization to inputs far longer than those seen in training. Our experiments on SALMONN and Qwen2-Audio confirm that Partial YaRN outperforms the original models across wide range of settings, and VLAT provides substantial performance improvement on long audio of unseen lengths. [^1] Code and dataset are available at: https://github.com/yophis/partial-yarn.
@@
Yuatyong Chaichana1,2thanks: Work done while interning at SCB 10X., Pittawat Taveekitworachai122footnotemark: 2, Warit Sirichotedumrong122footnotemark: 2, Potsawee Manakul122footnotemark: 2, Kunat Pipatanakul1thanks: Equal advising. 1SCB 10X, SCBX Group, 2Chulalongkorn University Correspondence: 6531337321@student.chula.ac.th
Figure 1: An illustration of Partial YaRN. The top plot displays low (blue) and high (red) frequency dimensions of RoPE. The bottom plot displays RoPE with Partial YaRN applied. In this example, the audio window of the low frequency dimensions are interpolated (stretched) by a factor of 1.75x to accommodate longer audio inputs, while the high frequency dimensions remain unaltered to preserve local distances and high frequency information (Peng et al., 2024). Within the Extended Audio Region, the magnitudes of every dimension is scaled to serve as a proxy for attention temperature scaling. Despite that the modality-decoupled scaling can create a minor discontinuity in the signal, information from other dimensions can help the model disambiguate this.
@@ 1 Introduction
The success of Large Language Models (LLMs) has spurred multimodal extensions, notably large audio-language models (LALMs) [^2] Also known by other names such as Speech-aware LLMs. that pair an audio encoder with a text backbone. By aligning audio and text into a shared representation space, LALMs can leverage the base LLM’s knowledge for complex audio understanding. However, practical use is constrained by short audio context windows: models such as SALMONN (Tang et al., 2024) and Qwen2-Audio (Chu et al., 2024) are typically trained on audio segments of 30s or less, and thus generalize poorly to longer inputs (Guo et al., 2025).
@@
This potential drawback motivates a more targeted strategy. We introduce Partial YaRN, a modality-decoupled, audio-only context extension method that modifies only the audio tokens’ positional encodings. This design enables a direct comparative study against the whole-context approaches, allowing us to investigate the tradeoff between preserving position encodings of the text modality and maintaining a globally uniform positional space.
Our experiments on both training-free and fine-tuned audio context extensions reveal that extension-based methods outperform original models. However, the best choice between whole-context and modality-decoupled extensions is often model-dependent without an universally superior solution. This suggests that to build truly robust long-audio-capable models, we must address the core problem of generalization during the training process itself. Therefore, we extend Partial YaRN into a novel fine-tuning technique called Virtual Longform Audio Training (VLAT). Acting as a positional augmentation technique, it simulates a wide range of audio lengths during fine-tuning, teaching the model to generalize beyond the lengths present in the training dataset. Through VLAT, we obtain substantially better generalization to longform audio of unseen lengths.
Our main contributions are as follows:
@@
  • We propose Virtual Longform Audio Training (VLAT), a training strategy that extends Partial YaRN into a positional augmentation to simulate diverse audio lengths during training, for improved long-context generalization.
  • We present a comparative study between existing context extension methods designed for unimodal LLMs and our proposed methods for LALMs, analyzing their performance tradeoff in both training-free and fine-tuning settings.
@@ 2.1 Rotary Positional Encoding
As Transformer (Vaswani et al., 2017) utilizes attention mechanism (Bahdanau et al., 2014), which is position-invariant, explicit positional information has to be provided for the models to differentiate tokens from different positions. Our study focuses on LALMs that use Rotary Positional Encoding (RoPE) (Su et al., 2024), a widely adopted positional encoding.
@@
$\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{k}_{n}$ $=(\boldsymbol{R}_{m,\boldsymbol{\theta}}\boldsymbol{q}_{m})^{\text{T}}(\boldsymbol{R}_{n,\boldsymbol{\theta}}\boldsymbol{k}_{n})$
$=\boldsymbol{q}_{m}^{\text{T}}\boldsymbol{R}_{n-m,\boldsymbol{\theta}}\boldsymbol{k}_{n}\,$,
@@ 2.2 Context Length Extension
Chen et al. (2023) proposes that instead of extrapolating RoPE further to the region unfamiliar for the trained model, interpolation is done inside the pretrained context window. Specifically, let $L$ be the original context length of the model, and $L^{\prime}$ be the target length to be extended to. Define $s=\frac{L^{\prime}}{L}$ as the extension factor. PI simply downscales the base rotational frequency by the extension factor: $\frac{\theta_{i}}{s}$. For example, if originally each positional step is rotated by $\theta_{i}=10^{\circ}$; to double the context length, the frequency (step size) can simply be halved to $5^{\circ}$. This way, the model will never have to attend to positions outside of its trained window, resulting in a more stable approach.
@@
$\operatorname{AS}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}\left(\frac{\boldsymbol{Q}^{T}\boldsymbol{K}}{t\sqrt{d_{k}}}\right)\boldsymbol{V},$
@@
Ding et al. (2024) proposes LongRoPE to account for non-uniformity across RoPE dimensions through a novel evolutionary search algorithm to determine dimension-wise interpolation factors. This method is further improved in LongRoPE2 (Shang et al., 2025).
@@ 2.3 Large Audio-language Models and Their Audio Contexts
One broadly adopted paradigm of large audio-language models (LALMs) is the unified input space (Tang et al., 2024; Chu et al., 2024; Ding et al., 2025; Ghosh et al., 2025a) which involves transforming audio inputs into the textual input space of a base LLM. By sharing the embedding space, the strong text capability and learned knowledge of the base LLM can be leveraged while simultaneously being augmented with audio understanding ability. There also exists other LALM paradigms such as cross attention (Kong et al., 2024; Ghosh et al., 2025b) where the audio and text modalities are fused through cross-attention modules. This work focuses on models under the unified-input-space.
As with text context window in LLMs, LALMs also do not generalize well to audio with lengths much longer than the window seen during training (e.g. 7s for Pengi (Deshmukh et al., 2023); 30s for Qwen2-Audio (Chu et al., 2024)).
@@ 3 Partial YaRN
To extend existing unified input LALMs to longer audio, we hypothesize that “LALMs already possess a sufficient general understanding of audio and text, the main bottleneck is instead their unfamiliarity to audio positions beyond the range seen during audio-text training. [^3] This differs from whole-context extension in unimodal LLMs, where the total sequence length exceeds the pre-trained context window, causing an out-of-distribution problem. Here, the total sequence (audio + text) usually still fits within the original window, making the core challenge one of adapting to an unfamiliar audio length and positions, rather than extrapolating to completely unseen positions. This hypothesis suggests a training-free solution: manipulating the backbone LLM’s positional encoding. The goal is to remap positions of a long audio input into the model’s familiar audio range.
First, we study the application of YaRN (Peng et al., 2024) to LALMs, as an example of whole-context extension methods. Since the audio window is a part of the total context window, performing whole-context extension extends the audio window as a byproduct. However, this alters positional information for all tokens, including text, risking the degradation of the base LLM, which was pretrained solely in text. Motivated by the aforementioned hypothesis, we conceive modality-decoupled extension, and propose Partial YaRN, an audio-only extension method designed to exclusively stretch the audio window. By leaving the text’s positional encodings unaltered, this approach aims to extend [^4] We use the terms extend and stretch interchangeably. audio context while preserving the base model’s text capability.
@@ 3.1 Methodology
Inspired by PI (Chen et al., 2023) and YaRN (Peng et al., 2024) for whole-context extension in unimodal LLMs, we adapt the interpolation technique to LALMs in modality-decoupled manner, applying it exclusively to the audio region of the base language models.
@@
where $\oplus$ denotes concatenation in the temporal axis. Note that while the final position index is denoted by $L$, the total token capacity now increases to $L+L^{\prime}_{\text{audio}}-L_{\text{audio}}$ to account for the stretched audio window with higher token frequency.
@@
We provide empirical validation and a more in-depth discussion of the two-group partitioning in Section 7.2 and Appendix A respectively. See Figure 1 for an example depiction of Partial YaRN.
@@ 4.1 Data
We source audio from the English subset of YODAS2 (Li et al., 2023). We segment audio samples into non-overlapping segments of 1, 2, 5, and 10 minutes. For each segment, we generate five multiple-choice question-answering (MCQA) pairs using Gemini 2.0 Flash (Team et al., 2023), with four choices each. We design the generation prompt to ensure that each question focuses on different portions of the audio, and that they collectively cover the entire audio segment. The test set of each audio duration (1, 2, 5, and 10 minutes) has 750 QA pairs.
@@ 4.3 Baseline Methods
These three baselines require no tuning as Vanilla and Whole PI don’t utilize any hyperparameter, and YaRN has predefined cutoffs and a closed-form formula for temperature. [^5] Unlike Whole YaRN for unimodal LLMs, where information per token is relatively consistent, compression of audio tokens varied significantly across different LALMs as mentioned. This likely influences the optimal hyperparameter choice considerably, as discussed in Appendix E.2.
@@
We also report performance of two proprietary models capable of long audio context: GPT-4o (OpenAI, 2024) and Gemini 2.0 Flash (Team et al., 2023), in order to provide a broader perspective of LALMs and to validate the quality of our dataset. [^6] The model versions are gpt-4o-mini-2024-07-18 and gemini-2.0-flash-001 respectively.
@@ 5 Main Results
In this section, we comparatively study the effect of whole-context and modality-decoupled context extensions on both the training-free and fine-tuning settings for audio segments up to 10mins.
@@ 5.1 Training-free Audio Context Extension
First, we extend the audio context length without training. We report the results in Table 4.1, showing that: no single method—whole-context or modality-decoupled—is universally superior. Instead, the performance varies between the different models and the degree of extension. More importantly, we observe that both models generally retain consistent performance up to 2mins, suggesting that their innate audio context lengths are longer than just 30s, possibly a result from multi-audio training.
@@
As mentioned in the beginning of this section that we can observe both model’s innate audio contexts to be around 2mins instead of just 30s, here we experiment with extending from this observed audio context of 2mins, to the target lengths. For both SALMONN and Qwen2-Audio, performance on the long audio improves substantially when the interpolation is anchored from the 2mins context instead of the original 30s. Notably on Qwen2-Audio, this strategy boosts Partial YaRN’s performance at 10mins from 28.53% to a 48.00%, an absolute improvement of almost 20%, yielding a final performance 26% higher than the Vanilla baseline. This finding suggests that determining the extension by observing the vanilla performance may be a more critical factor than the choice between whole-context or modality-decoupled methods, especially on large extension ratios.
@@ 5.2 Fine-tuning for Audio Context Extension
The results are presented in Table 3, showing that fine-tuning with a context extension method dramatically outperforms the vanilla fine-tuning baseline, especially at longer audio lengths. In the qkvo setting, both Whole YaRN (83.47%) and Partial PI (83.07%) achieve an absolute performance gain of around 19% over the Vanilla baseline (64.93%) at 10mins. This highlights the benefit of integrating a context extension method into the fine-tuning process for better context extension.
Comparing the whole-context and modality-decoupled approaches, we find that their performance is generally competitive. While Whole YaRN triumps when only adapting the query projection weights, suggesting its ease of adaptation. Their performance converges under standard LoRA practice which adapts qkvo. This result confirms that both whole-context and audio-only methods are highly effective strategies for fine-tuning.
@@
Table 3: Fine-tuned audio context extension performance (Accuracy). We fine-tune Qwen2-Audio on YODAS2-MCQA, under two LoRA settings: adapting only q, and adapting qkvo. Each model is fine-tuned and evaluated using the same extension method.
@@ 6 Virtual Longform Audio Training: Generalizing Beyond the Training Data
Our previous results show that while the training-free methods are beneficial, their effectiveness is often model and length dependent. Based on this, we turn to address the core generalization problem during the training process, and propose to fine-tune LALMs with Partial YaRN repurposed as a positional augmentation technique we call “Virtual Longform Audio Training” (VLAT).